# Irredundant k–Fold Cross–Validation

### v1

@@ 5 Conclusions
The proposed Irredundant $k$-fold cross–validation retains the conceptual simplicity of classical $k$–fold cross–validation while removing a single but consequential source of redundancy: no instance ever appears in more than one training fold. This simple constraint confers several main advantages.
@@
[…]
The irredundant design of $\mathsf{IkF}$ impacts remarkably on the computational cost of the validation ($\approx$2.9×2.9$\times$2.9 × speed–up), which might be highly significant for complex models such as those based on neural architectures (e.g., convolutional neural networks, generative adversarial networks, transformers, etc.).

### v2

@@ 5 Conclusions
The proposed Irredundant $k$-fold cross–validation retains the conceptual simplicity of classical kk–fold cross–validation while removing a single but consequential source of redundancy: no instance ever appears in more than one training fold. This simple constraint confers several main advantages.
@@
Although each model is trained on $n/k$ rather than (k1)n/k(k-1)n/k samples—introducing a slightly conservative bias—the empirical gap in accuracy and F–score is marginal ($\approx$2.7 and $\approx$4.8, respectively). In settings where under–fitted behavior is preferable to over–optimism (e.g., medical AI validation), the $\mathsf{IkF}$ validation properties can be an advantage.
The irredundant design of $\mathsf{IkF}$ impacts remarkably on the computational cost of the validation ($\approx$$2.9\times$ speed–up), which might be highly significant for complex models such as those based on neural architectures (e.g., convolutional neural networks, generative adversarial networks, transformers, etc.).